OpenStudy (anonymous):

y = 8x^2 + 2x + 2/sqrootx

OpenStudy (lgbasallote):

and....?

OpenStudy (anonymous):

\[\frac{ \sqrt{2}(16x+2)-8x^{2 ^{^{}}}+2x+2(\frac{ 1 }{ 2 }x ^{-\frac{ 1 }{ 2 }}) }{ \sqrt{x}^{2} }\]

OpenStudy (anonymous):

Differentiate the function.

OpenStudy (anonymous):

there is a x in the sqroot not 2